OpenStudy (anonymous):

solve by completing the square,what value should be added to each side of the equation? x^2 + 14x = -7

OpenStudy (anonymous):

1/2 the middle term squared so 14 is the middle term number 14/2 = 7 7^2 = 49

OpenStudy (anonymous):

then notice that x^2 + 14x + 49 is factorable

OpenStudy (whpalmer4):

Don't even have to notice that — if you've added the square of half the coefficient of the x term, then the left hand side becomes (x+14/2)^2 or (x+7)^2

OpenStudy (whpalmer4):

though yes, hopefully you do notice it :-)

OpenStudy (anonymous):

i was only pointing it out, because if you did the process incorrectly, your result would not be factorable since it was factorable, must be the right process lol

OpenStudy (anonymous):

so if I add (x+7)^2 to each side I can square it?

OpenStudy (anonymous):

noo you dont need to make a binomial look at it this way x^2 + 14x + C you want to find some number C, so that you can factor the whole thing

OpenStudy (anonymous):

to do that, just create the value of C from the other numbers namely, the coefficient of your middle term, which is 14

OpenStudy (anonymous):

if you divide 14 by 2 (take half of it), then square that number, you will create a number which can be factored

OpenStudy (anonymous):

the reason you take half, is because you are going to factor into 2 terms, and those two terms need to add up to 14 if you take half of 14, you get 7 and 7, which adds up perfectly to 14 the reason you square 7, is because you are factoring, and need a number which has a perfect square root (49, in this case)

OpenStudy (whpalmer4):

mmm, no, you're just rewriting the left hand side as a perfect square. \[(x+14/2)^2 = (x+7)^2 = x^2+14x+49\]You had to add something to both sides so that the left side fit into that form. Now you have\[(x+7)^2 = -7 + 7^2\]\[(x+7)^2=42\]You can take the square root of both sides and find \(x\): \[\sqrt{(x+7)^2} = \sqrt{42}\]\[\pm(x+7) = \sqrt{42}\]You have to solve with both the positive and negative square roots to get both solutions, so that is shorthand for \[(x+7) = \sqrt{42}\]\[-(x+7) = \sqrt{42}\]